# h FR-PHENO-2025-02 A general approach to quantum integration of cross sections in high-energy physics

### v1

@@ 4 Example applications
[…]
Figure 8: Feynman diagram for decay of a tau fermion [see Eq. 29].
@@
[…]
@@ 4.1 One-dimensional integration
To demonstrate the methodology, for simplicity, we start by considering only the numerator of Eq. 30. By neglecting the constant factors and integrating only over the two invariants $s_{1}$ and $s_{1}$, one obtains the following integral
@@
which has analytical value $\sigma=\frac{s^{4}}{12}-m^{2}_{\tau}\frac{s^{3}}{6}$
@@
where $C(s_{1},s_{2})=1$, if $s_{1}+s_{2}<s$, and $C(s_{1},s_{2})=0$, otherwise. If we compare this to the general form of the expectation calculated using the QMCI engine in Eq. 7, and identify $s_{1}=x$ and $s_{2}=y$ (this identification will remain for all subsequent examples), we observe that there is freedom in the integrand separation, in terms of what to define as the probability distribution, $f_{S_{1}S_{2}}(s_{1},s_{2})$, and the function applied $g(s_{1},s_{2})$. In this case, for simplicity, we choose to set $f_{S_{1}S_{2}}(s_{1},s_{2})=U(s_{1})U(s_{2})$, where $U(.)$ is the uniform distribution, and $g(s_{1},s_{2})=s_{1}^{2}$ or $g(s_{1},s_{2})=s_{1}$. [^14] Note that this choice does not leverage the capabilities of the QMCI engine in terms of decomposing the integrand into building blocks, as discussed previously—however, that is not the purpose of this initial example.
@@
As discussed previously in Section 2.3.2, the QMCI engine contains powerful, in-built functionality for performing such a calculation; the cut function can be straightforwardly implemented using thresholding operations, whilst the engine contains efficient methods for calculating both $g(s_{1},s_{2})=s_{1}^{2}$ and $g(s_{1},s_{2})=s_{1}$. The QMCI engine can thus be used to build efficient, low depth circuits that give estimates of both expectation values on the RHS of Eq. 33.
@@
For this example, and the ones described later in Sections 4.2 and 4.3, we considered three different levels of expected precision, corresponding to upper bounds on the expected RMSE of the final estimator of the order $10\%$, $1\%$, and $0.1\%$, [^15] The per-mille precision represents the typical accuracy of classical MCI calculation in HEP. respectively (not accounting for the systematic errors related to the state preparation discussed in Section 3). For illustration purposes, we run noiseless simulations of the circuits for the case of $10\%$ precision.
[…]
@@
In order to demonstrate the validity of the approach, we ran noiseless simulations using the QMCI engine to numerically estimate the RHS of Eq. 33, which we denote as $\hat{\sigma}$, $100$ different times, based on an expected precision of $10\%$. The numerical experiments, and the ones described later in Sections 4.2 and 4.3, were carried out using Qulac’s noiseless state-vector simulator [66]. The QAE algorithm used in all cases was an in-built implementation of the maximum-likelihood QAE algorithm [67], that is optimised to make use of all available samples in order to maximise performance.
@@
Figure 9 gives a histogram of the error, $\hat{\sigma}-\sigma$, where we can see that the results are as expected, with nearly all (except for a single outlier) values falling within an expected precision of $10\%$.
@@
Figure 9: Histogram of the error for $100$ runs of QMCI estimating the RHS of Eq. 33, for an expected RMSE upper-bound precision of $10\%$.
@@
Using the in-built resource mode of the QMCI engine (described in Section 2.3.2), we also calculated the resources required to perform this calculation to the three different levels of expected precision. Table 2 gives these resources for both NISQ and fault-tolerant compilation. We note that the number of qubits required is in general small [$\mathcal{O}(10)$], and in particular is well within the reach of current quantum hardware (this is assuming that the systematic errors based on using $5$-qubit distributions do not excessively affect calculations—something that is not observed for this case for $10\%$ precision). However, the number of operations required for NISQ compilation almost certainly means that logical qubits would be required; considering the largest circuit for an expected precision of $10\%$ has $1.80\times 10^{6}$ CX gates, then a simple rule-of-thumb calculation suggests one would need a machine with a two-qubit gate fidelity of approximately $1-1/(1.80\times 10^{6})=99.99994\%$—well out of reach of current quantum hardware. We note that the number of T-gate operations for fault-tolerant execution are also high; for example, for a $0.1\%$ precision, as would be required to compete with current classical MCI methods, the largest circuit requires $1.71\times 10^{9}$ T gates. However, it is worth noting these gate counts are likely to reduce significantly in future as more research into optimised synthesis is carried out (see Ref. [32], Sec. 10.3 for a detailed discussion of this topic), and, therefore, these values should be regarded as loose upper bounds, and treated with a degree of uncertainty.
@@
Table 2: Resources required to estimate the RHS of Eq. 33 using QMCI for various expected precisions, for both NISQ and fault-tolerant resource mode.
@@
Number in largest circuit $1.80\times 10^{6}$ $1.14\times 10^{7}$ $1.69\times 10^{8}$
@@ 4.2 Separable two-dimensional integration
We now move on to the expression for the actual integration of the tau decay given in Eq. 30, by also considering the denominator (i.e., the propagator). We again omit the integration over the angles. The integral then becomes
@@
The analytical solution to this integral is given in Appendix A.
@@
where the cut function, $C(s_{1},s_{2})$, is the same as previously. Then, recalling the general form of an expectation calculated using the QMCI engine in Eq. 7, and following the discussion in Section 2.4, we see that each individual integral has the canonical form, in terms of products of building blocks, that we require for performing a generic cross-section calculation using the QMCI engine. We can identify the probability distribution, $f_{s_{1}s_{2}}$, as a product of univariate BW distributions [or rather, in this case, the product of an uniform distribution for the dimension $s_{1}$, and an univariate W-boson BW distribution for the dimension $s_{2}$, i.e., $f_{s_{1}s_{2}}(s_{1},s_{2})=U(s_{1})\text{BW}_{\text{W}}(s_{2})$]. Then, the functions applied, $g(.)$, are just products of moments of the integration variables [or rather in this case simply the univariate products $g(s_{1},s_{2})=s_{1}^{2}$ or $g(s_{1},s_{2})=s_{1}$].
@@
For this example, for demonstration purposes, it is important that we integrate across the range of the BW peak, and therefore we set $\sqrt{s_{max}}=100\,\text{GeV}$. [^17] Note that this calculation does not correspond to a physical tau decay (or indeed a physical process). The analytical value in this case is $\sigma=3.162\times 10^{8}$. As discussed in Section 3, we make use of a trained PQC with $n=6$ (specifically the optimal circuit given in Table 1) to prepare the BW distribution for the W-boson propagator used to model the probability distributions for dimension $s_{2}$, and use the QMCI engine’s state-preparation library to load a $n=6$ circuit preparing the uniform distribution used to model the probability distribution for dimension $s_{1}$. We again construct $f_{s_{1}s_{2}}(s_{1},s_{2})$ by combining the circuits.
@@
In order to demonstrate the validity of our generic approach, we carried out numerical experiments by running noiseless simulations to numerically estimate the RHS of Section 4.2, $20$ different times, for an expected precision of $10\%$.
@@
Figure 10 gives a histogram of the error, $\hat{\sigma}-\sigma$, where we again see all values within the expected precision of $10\%$. However, in this case there is a clear bias in the results (as they are not centered around 00). A bias is to be expected given the discussion of systematic errors regarding the state preparation of BW distributions given in Section 3—a source of error that was notably not applicable to the previous example. In this case, the observed systematic bias is approximately an order of magnitude smaller than the required precision, and therefore negligible. However, for higher levels of precision—such as for $1\%$ and $0.1\%$—one would require state-preparation circuits with smaller corresponding systematic errors, in order to still obtain accurate results.
@@
Figure 10: Histogram of the error for $20$ runs of QMCI estimating the RHS of Section 4.2, for an expected RMSE upper-bound precision of $10\%$.
@@
We used the resource mode to calculate the resources for the three expected levels of precision. Table 3 gives the resources for NISQ and fault-tolerant compilation. The results are similar to those found in Section 4.1, and therefore a similar analysis to that described there applies. However, in this case it is worth noting that due to the more complicated circuit used to prepare the BW distribution for the $s_{2}$ dimension, then the qubit counts are slightly larger, and the gate counts are also approximately an order of magnitude larger.
@@
Table 3: Resources required to estimate the RHS of Section 4.2 using QMCI for various expected precisions, for both NISQ and fault-tolerant resource mode.
@@ 4.3 Non-separable two-dimensional integration
Finally, in order to increase the complexity of the problem and mimic the more general case of Eq. 4, we compute the following
[…]
which amounts in practice to, in addition to calculating the same terms as in Section 4.2 (multiplied by the constant, $s$), calculating the following, non-separable, two-dimensional integral
[…]
The analytical solution to this additional integral is also provided in Appendix A. The analytical value of Eq. 36 is then $\sigma=3.179\times 10^{12}$.
For this example, the setup is the same as in Section 4.2. In order to demonstrate the validity of our approach, we carried out numerical experiments by running noiseless simulations to numerically estimate the RHS of Eq. 36, 4 times, for an expected precision of $10\%$.
Figure 11 gives a histogram of the error, $\hat{\sigma}-\sigma$. The results are broadly similar to those in Section 4.2, with all values within the expected precision, and with a bias again observed (this is not surprising, as the calculations only differ by the additional term, which is subdominant).
Figure 11: Histogram of the error for 4 runs of QMCI estimating the RHS of Eq. 36 for an expected RMSE upper-bound precision of $10\%$.
We analysed the resources required to run the circuits at the nominal precisions using the resource mode. Table 4 gives these for NISQ and fault-tolerant compilation. Once again the results are similar to the previous sections, and therefore a similar analysis again applies. In this case, as compared to the example in Section 4.2—which itself required greater resources than the example in Section 4.1—both the qubit counts and number of gates are larger, by approximately an additional order of magnitude in the latter case. This increase is due to the extra 2D integration that is performed (i.e., calculating $\mathbb{E}[s_{1}s_{2}]$), which as discussed in Section 2.3.1 requires greater resources than when calculating univariate expectations. Thus, we see that as the dimensionality or complexity of the integrals increases, so too do the required resources.
@@
Table 4: Resources required to estimate the RHS of Eq. 36 using QMCI for various expected precisions, for both NISQ and fault-tolerant resource mode.

### v2

@@ 4 Example applications
where the invariants, $s_{1}$ and $s_{2}$, are equal to $(p_{1}+p_{3})^{2}$ and $(p_{2}+p_{3})^{2}$, respectively. The electroweak coupling is denoted by $\alpha$, while the weak mixing angle is $\theta_{\rm w}$. The mass and width of the W bosons are denoted by $m_{\text{W}}$ and $\Gamma_{\text{W}}$, respectively, while $M_{\tau}$ is the mass of the tau lepton. [^17] It is worth noting that the expression for the matrix element in Eq. 30 assumes a massless electron, which is a common approximation in HEP. This is because the mass of the electron ($0.5\,\text{MeV}$) is negligible with respect to a collision energy ($13.6\,\text{TeV}$ for the LHC, for example). It is worth pointing out that the expression in Eq.˜30 contains a W-boson propagator term [see Eq.˜3]. This originates from the fact that the electron and anti-electron neutrino are produced through the decay of an intermediate W boson, as depicted in the Feynman diagram in Fig.˜8.
Figure 8: Feynman diagram for decay of a tau fermion [see Eq.˜29].
@@
$={}$ $\frac{1}{32s}\mathrm{d}s_{1}\mathrm{d}s_{2}\mathrm{d}\Omega_{1}\mathrm{d}\phi_{3},$ (31)
@@ 4.1 One-dimensional integration
To demonstrate the methodology, for simplicity, we start by considering only the numerator of Eq.˜30. By neglecting the constant factors and integrating only over the two invariants $s_{1}$ and $s_{1}$, one obtains the following integral
@@
which has analytical value $\sigma=\frac{s^{4}}{12}-M_{\tau}^{2}\frac{s^{3}}{6}$
@@
where $C(s_{1},s_{2})=1$, if $s_{1}+s_{2}<s$, and $C(s_{1},s_{2})=0$, otherwise. If we compare this to the general form of the expectation calculated using the QMCI engine in Eq.˜7, and identify $s_{1}=x$ and $s_{2}=y$ (this identification will remain for all subsequent examples), we observe that there is freedom in the integrand separation, in terms of what to define as the probability distribution, $f_{S_{1}S_{2}}(s_{1},s_{2})$, and the function applied $g(s_{1},s_{2})$. In this case, for simplicity, we choose to set $f_{S_{1}S_{2}}(s_{1},s_{2})=U(s_{1})U(s_{2})$, where $U(.)$ is the uniform distribution, and $g(s_{1},s_{2})=s_{1}^{2}$ or $g(s_{1},s_{2})=s_{1}$. [^18] Note that this choice does not leverage the capabilities of the QMCI engine in terms of decomposing the integrand into building blocks, as discussed previously—however, that is not the purpose of this initial example.
@@
As discussed previously in Section˜2.3.2, the QMCI engine contains powerful, in-built functionality for performing such a calculation; the cut function can be straightforwardly implemented using thresholding operations, whilst the engine contains efficient methods for calculating both $g(s_{1},s_{2})=s_{1}^{2}$ and $g(s_{1},s_{2})=s_{1}$. The QMCI engine can thus be used to build efficient, low depth circuits that give estimates of both expectation values on the RHS of Eq.˜33.
@@
For this example, and the ones described later in Sections˜4.2 and 4.3, we considered three different levels of expected precision, corresponding to upper bounds on the expected RMSE of the final estimator of the order $0\%, 11\%$, and $0.1\%$, [^19] The per-mille precision represents the typical accuracy of classical MCI calculation in HEP. respectively (not accounting for the systematic errors related to the state preparation discussed in Section˜3). For illustration purposes, we run noiseless simulations of the circuits for the case of $10\%$ precision.
We considered a decay-like process where $\sqrt{s_{max}}=M_{\tau}=1.776\,\text{GeV}$, giving an analytical value $\sigma=-8.248$. [^20] Given that this quantity is a proxy of a cross section, it is not physical, and thus not necessarily positive. We set $n=5$ for each of the dimensions, $s_{1}$ and $s_{2}$, respectively. This results in a final QMCI circuit containing $23$ qubits in total. The circuits to prepare the uniform distributions used to model the probability distributions for each dimension were loaded directly in the QMCI engine using the in-built state-preparation library, and these then combined to construct the bivariate probability distribution, $f_{S_{1}S_{2}}(s_{1},s_{2})$.
@@
In order to demonstrate the validity of the approach, we ran noiseless simulations using the QMCI engine to numerically estimate the RHS of Eq.˜33, which we denote as $\hat{\sigma}$, $100$ different times, based on an expected precision of $10\%$. The numerical experiments, and the ones described later in Sections˜4.2 and 4.3, were carried out using Qulacs’ noiseless state-vector simulator [67]. The QAE algorithm used in all cases was an in-built implementation of the maximum-likelihood QAE algorithm [68], that is optimised to be able to run for any given number of samples, thereby maximising performance.
@@
Figure˜9 gives a histogram of the error, $\hat{\sigma}-\sigma$, where we can see that the results are as expected, with nearly all (except for a single outlier) values falling within an expected precision of $10\%$.
@@
Figure 9: Histogram of the error for $100$ runs of QMCI estimating the RHS of Eq.˜33, for an expected RMSE upper-bound precision of $10\%$.
@@
Using the in-built resource mode of the QMCI engine (described in Section˜2.3.2), we also calculated the resources required to perform this calculation to the three different levels of expected precision. Table˜2 gives these resources for both NISQ and fault-tolerant compilation. We note that the number of qubits required is in general small $\mathcal{O}(10)$], and in particular is well within the reach of current quantum hardware (this is assuming that the systematic errors based on using $5$-qubit circuits to prepare the uniform distributions do not excessively affect calculations—something that is not observed for this case for $10\%$ precision). However, the number of operations required for NISQ compilation almost certainly means that logical qubits would be required; considering the largest circuit for an expected precision of $10\%$ has $1.80\times 10^{6}$ CX gates, then a simple rule-of-thumb calculation suggests one would need a machine with a two-qubit gate fidelity of approximately $1-1/(1.80\times 10^{6})=99.99994\%$—well out of reach of current quantum hardware. Given the substantial two-qubit gate counts, even the application of targeted error mitigation strategies—such as noise-aware QAE techniques [69]—is unlikely to suffice. There is only so much error mitigation can do in the presence of such high noise levels, and the accumulation of errors is likely to compromise the accuracy of the results, regardless. We note that the number of T-gate operations for fault-tolerant execution are also high; for example, for a $0.1\%$ precision, as would be required to compete with current classical MCI methods, the largest circuit requires $1.71\times 10^{9}$ T gates. However, it is worth noting these gate counts are likely to reduce significantly in future as more research into optimised synthesis is carried out (see Section 10.3 of Ref. [32] for a detailed discussion of this topic), and, therefore, these values should be regarded as loose upper bounds, and treated with a degree of uncertainty.
@@
Table 2: Resources required to estimate the RHS of Eq.˜33 using QMCI for various expected precisions, for both NISQ and fault-tolerant resource mode.
@@
Number in largest circuit $80\times 10^{6} 1.1$$14\times 10^{7}$ $1.69\times 10^{8}$
@@ 4.2 Separable two-dimensional integration
We now move on to the expression for the actual integration of the tau decay given in Eq.˜30, by also considering the denominator (i.e., the propagator). We again omit the integration over the angles. The integral then becomes
@@
The analytical solution to this integral is given in Appendix˜A.
@@
where the cut function, $C(s_{1},s_{2})$, is the same as previously. Then, recalling the general form of an expectation calculated using the QMCI engine in Eq.˜7, and following the discussion in Section˜2.4, we see that each individual integral has the canonical form, in terms of products of building blocks, that we require for performing a generic cross-section calculation using the QMCI engine. We can identify the probability distribution, $f_{s_{1}s_{2}}$, as a product of univariate BW distributions [or rather, in this case, the product of an uniform distribution for the dimension $s_{1}$, and an univariate W-boson BW distribution for the dimension $s_{2}$, i.e. $f_{s_{1}s_{2}}(s_{1},s_{2})=U(s_{1})\text{BW}_{\text{W}}(s_{2})$]. Then, the functions applied, $g(.)$, are just products of moments of the integration variables [or rather in this case simply the univariate products $g(s_{1},s_{2})=s_{1}^{2}$ or $g(s_{1},s_{2})=s_{1}$].
@@
For this example, for demonstration purposes, it is important that we integrate across the range of the BW peak, and therefore we set $\sqrt{s_{max}}=100\,\text{GeV}$. [^21] Note that this calculation does not correspond to a physical tau decay (or indeed a physical process). The analytical value in this case is $\sigma=3.162\times 10^{8}$. As discussed in Section˜3, we make use of a trained PQC with $n=6$ (specifically the optimal circuit given in Table˜1) to prepare the BW distribution for the W-boson propagator used to model the probability distributions for dimension $s_{2}$, and use the QMCI engine’s state-preparation library to load a $n=6$ circuit preparing the uniform distribution used to model the probability distribution for dimension $s_{1}$. This results in a QMCI circuit containing $27$ qubits in total. We again construct $f_{s_{1}s_{2}}(s_{1},s_{2})$ by combining the circuits.
@@
In order to demonstrate the validity of our generic approach, we carried out numerical experiments by running noiseless simulations to numerically estimate the RHS of Section˜4.2, $24$ different times, for an expected precision of $10\%$.
@@
Figure˜10 gives a histogram of the error, $\hat{\sigma}-\sigma$, where we again see all values within the expected precision of $10\%$. However, in this case there appears to be a bias in the results (as they are not centred around 0). Some bias is likely expected given the discussion of systematic errors regarding the state preparation of BW distributions given in Section˜3—a source of error that was notably not applicable to the previous example. In this case, the observed bias is approximately an order of magnitude smaller than the required precision, and is therefore negligible. However, for higher levels of precision—such as for $1\%$ and $0.1\%$—one would likely require state-preparation circuits with smaller corresponding systematic errors, in order to still obtain accurate results. In particular, a quantitative analysis of the effect of such systematic errors on the final estimation error is needed to be able to fully characterise these QMCI calculations. However, such a detailed analysis would require a significant extension to the error analyses detailed in Section˜3 (which are derived from Ref. [33]). This is because, for this particular example, systematic errors from the state preparation only indirectly affect the final estimation error via the thresholding condition; the $s_{2}$ dimension is not directly integrated over, but is summed with the $s_{1}$ dimension to form a new dimension, and this new dimension used for the threshold condition, thereby indirectly propagating the error in the state preparation to the final estimation. An extension of the error analyses to be able to provide a quantitative treatment for such cases is well beyond the scope of this article, constituting an independent research direction, and is therefore left for future work.
@@
Figure 10: Histogram of the error for $24$ runs of QMCI estimating the RHS of Section˜4.2, for an expected RMSE upper-bound precision of $10\%$.
@@
We used the resource mode to calculate the resources for the three expected levels of precision. Table˜3 gives the resources for NISQ and fault-tolerant compilation. The results are similar to those found in Section˜4.1, and therefore a similar analysis to that described there applies. However, in this case it is worth noting that due to the more complicated circuit used to prepare the BW distribution for the $s_{2}$ dimension, then the qubit counts are slightly larger, and the gate counts are also approximately an order of magnitude larger.
@@
Table 3: Resources required to estimate the RHS of Section˜4.2 using QMCI for various expected precisions, for both NISQ and fault-tolerant resource mode.
@@ 4.3 Non-separable two-dimensional integration
Finally, in order to increase the complexity of the problem and mimic the more general case of Eq.˜4, we compute the following
$\sigma\propto{}$ $\int^{s}_{0}\int^{s-s_{2}}_{0}ds_{1}ds_{2}\frac{s_{1}^{2}s-s_{1}M_{\tau}^{2}s+s_{1}M_{\tau}^{2}s_{2}}{(s_{2}-M^{2}_{\text{W}})^{2}+(M_{\text{W}}\Gamma_{\text{W}})^{2}},$ (36)
which amounts in practice to, in addition to calculating the same terms as in Section˜4.2 (multiplied by the constant, $s$), calculating the following, non-separable, two-dimensional integral
$I_{1}={}$ $\int^{s}_{0}\int^{s-s_{2}}_{0}ds_{1}ds_{2}\frac{s_{1}M_{\tau}^{2}s_{2}}{(s_{2}-M^{2}_{\text{W}})^{2}+(M_{\text{W}}\Gamma_{\text{W}})^{2}}.$ (37)
The analytical solution to this additional integral is also provided in Appendix˜A. The analytical value of Eq.˜36 is then $\sigma=3.179\times 10^{12}$.
For this example, the setup is the same as in Section˜4.2. In order to demonstrate the validity of our approach, we carried out numerical experiments by running noiseless simulations to numerically estimate the RHS of Eq.˜36, $6$ times, for an expected precision of $10\%$.
Figure˜11 gives a histogram of the error, $\hat{\sigma}-\sigma$. The results are broadly similar to those in Section˜4.2, with all values within the expected precision, and with a bias again observed (this is not surprising, as the calculations only differ by the additional term, which is subdominant).
Figure 11: Histogram of the error for 66 runs of QMCI estimating the RHS of Eq.˜36 for an expected RMSE upper-bound precision of $10\%$.
We analysed the resources required to run the circuits at the nominal precisions using the resource mode. Table˜4 gives these for NISQ and fault-tolerant compilation. Once again the results are similar to the previous sections, and therefore a similar analysis again applies. In this case, as compared to the example in Section˜4.2—which itself required greater resources than the example in Section˜4.1—both the qubit counts and number of gates are larger, by approximately an additional order of magnitude in the latter case. This increase is due to the extra 2D integration that is performed (i.e. calculating $\mathbb{E}[s_{1}s_{2}]$), which as discussed in Section˜2.3.1 requires greater resources than when calculating univariate expectations. Thus, we see that as the dimensionality or complexity of the integrals increases, so too do the required resources.
@@
Table 4: Resources required to estimate the RHS of Eq.˜36 using QMCI for various expected precisions, for both NISQ and fault-tolerant resource mode.